Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Combinatrics

Expert replies
by Stockmoose16 » Thu Sep 11, 2008 5:04 pm
A box contains one dozen donuts. Four of the donuts are chocolate, four are glazed, and four are jelly. If two donuts are randomly selected from the box, one after the other, what is the probability that both will be jelly donuts?

1/11
1/9
1/3
2/3
8/9

I'm know to solve this problem using probability, you could just do 4/12*3/11 =1/11.

However, when I try to solve using combinatrics, I'm wondering if there's 2 different ways to do it:

VERSION #1
Number of total possibilities 12C2 = 66
Number of ways to pick 2 Jelly donuts: 4C2= 6

6/66=1/11

OR

VERSION #2
Number of total possibilities: 12C1 * 11C1= 12*11= 132
Number of ways to pick 2 Jelly donuts= 4C1 * 3C1= 4*3=12

12/132=1/11

Are both these combinatrics methods viable? If so, how come they both work, even though version #2 involves multiplication and version #1 does not?
Join the discussion
Source: — Problem Solving |

Re: Combinatrics

by canuckclint » Thu Sep 11, 2008 8:31 pm
Stockmoose16 wrote:A box contains one dozen donuts. Four of the donuts are chocolate, four are glazed, and four are jelly. If two donuts are randomly selected from the box, one after the other, what is the probability that both will be jelly donuts?

1/11
1/9
1/3
2/3
8/9

I'm know to solve this problem using probability, you could just do 4/12*3/11 =1/11.

However, when I try to solve using combinatrics, I'm wondering if there's 2 different ways to do it:

VERSION #1
Number of total possibilities 12C2 = 66
Number of ways to pick 2 Jelly donuts: 4C2= 6

6/66=1/11

OR

VERSION #2
Number of total possibilities: 12C1 * 11C1= 12*11= 132
Number of ways to pick 2 Jelly donuts= 4C1 * 3C1= 4*3=12

12/132=1/11

Are both these combinatrics methods viable? If so, how come they both work, even though version #2 involves multiplication and version #1 does not?
Yes you are absolutely right. The second way instead of having c 1.
Just have probability of picking 1st jelly donut * 2nd jelly d onut
4/12 + 3/11 = 1/11, possibly version #2.5
Join the discussion

Re: Combinatrics

by canuckclint » Thu Sep 11, 2008 8:32 pm
canuckclint wrote:
Stockmoose16 wrote:A box contains one dozen donuts. Four of the donuts are chocolate, four are glazed, and four are jelly. If two donuts are randomly selected from the box, one after the other, what is the probability that both will be jelly donuts?

1/11
1/9
1/3
2/3
8/9

I'm know to solve this problem using probability, you could just do 4/12*3/11 =1/11.

However, when I try to solve using combinatrics, I'm wondering if there's 2 different ways to do it:

VERSION #1
Number of total possibilities 12C2 = 66
Number of ways to pick 2 Jelly donuts: 4C2= 6

6/66=1/11

OR

VERSION #2
Number of total possibilities: 12C1 * 11C1= 12*11= 132
Number of ways to pick 2 Jelly donuts= 4C1 * 3C1= 4*3=12

12/132=1/11

Are both these combinatrics methods viable? If so, how come they both work, even though version #2 involves multiplication and version #1 does not?
Yes you are absolutely right. The second way instead of having c 1.
Just have probability of picking 1st jelly donut * picking 2nd jelly donut
4/12 * 3/11 = 1/11, possibly version #2.5

Version 1 is not a good idea because (12 c 2) is very hard to compute by hand.
Join the discussion